# The Spectrum of Gravitational Waves from Annihilating Domain Walls

### v1

@@ Acknowledgements.
The authors thankfully acknowledge the use of computational resources from ICCUB Nyx cluster and the LXPLUS cluster at CERN. The work of F.T. is supported by a Beatriu de Pinós fellowship 2022-BP-00063 from the Ministry of Research of Catalonia. The work of F.R. is supported by the grant RYC2021-031105-I from the Ministerio de Ciencia e Innovación (Spain). The work of A.N. and F.T. is supported by the grants PID2019-108122GB-C32 from the Spanish Ministry of Science and Innovation, Unit of Excellence Maria de Maeztu 2020-2023 of ICCUB (CEX2019- 000918-M) and AGAUR 2021 SGR 00872. We thank Jorge Baeza-Ballesteros for his valuable assistance in producing snapshots of the lattice. We acknowledge fruitful interactions during completion of this work at the VIII Cosmological Olentzero Workshop in Bilbao, and thank the UPV for the warm hospitality.
@@ B.1 Initial conditions
We have evolved the fields in a radiation-dominated background $a(\eta)=a_{i}+H_{i}\eta$, with $a_{i}$ and $H_{i}$ the scale factor and Hubble parameter at the initial simulation time respectively. At that time we set the homogeneous mode of the inflaton amplitude to zero. Over it, we impose a white noise spectrum of Gaussian fluctuations in momentum space below a certain cutoff $k<k_{*}$. The resulting variance in position space is given by
@@
with $k_{*}$ a ultraviolet comoving cutoff. The procedure by which this spectrum is generated is explained in Sec. 7.1 of [28]. In our simulations we have fixed $k_{*}=2m$, but we have checked that the dynamics of the DW network are quite insensitive on the specific choice of $k_{*}$.
@@ B.3 Fourth-order accurate spatial derivatives
B.4 Spectrum dependence on lattice volume and $\eta_{\Delta V}$
[…]
On the other hand, in the simulations presented in Sec. III we have fixed the coupling constant of the bias term so that $\eta_{\Delta V}=14$. In order to assess the robustness of our spectrum parametrization under changes of the bias size, in Fig. 19 we compare the final GW spectrum obtained with one realization of the case $\eta_{\Delta V}=14$, with another one of the case $\eta_{\Delta V}=17$. We observe that both spectra are remarkably similar in the region $x<50$.
B.5 Comparison with previous results
In Fig. 19 we compare our prediction for the shape of the GW spectrum in the biased $\phi^{4}$ case with a time-independent bias, with the one obtained in Ref. [33] for a time-dependent one (see the top panel of Fig. 3 of that work). Both spectra agree quite well in the position of the peak and the slope of the UV tail in the region $0.3\lesssim k\lesssim 5$. On the other hand, the IR tail observed in [33] is steeper than ours. Both spectra also show a plateau at high frequencies, with a difference of a factor two in its comoving momentum.

### v3

@@ Acknowledgements.
The authors thankfully acknowledge the use of computational resources from the Nyx cluster at ICCUB, the LXPLUS cluster at CERN, and Finisterrae III at CESGA. The work of F.T. is supported by a Beatriu de Pinós fellowship 2022-BP-00063 from the Ministry of Research of Catalonia. The work of F.R. is supported by the grant RYC2021-031105-I from the Ministerio de Ciencia e Innovación (Spain). The work of A.N. and F.T. is supported by the grants PID2022-137268NB-C52 from the Spanish Ministry of Science and Innovation, Unit of Excellence Maria de Maeztu 2024-2027 of ICCUB (CEX2024-001451-M) and AGAUR 2021 SGR 00872. We thank Jorge Baeza-Ballesteros for his valuable assistance in producing snapshots of the lattice. We acknowledge fruitful interactions during completion of this work at the VIII Cosmological Olentzero Workshop in Bilbao, and thank the UPV for the warm hospitality.
@@ Appendix A Lattice simulations for the unbiased $\phi^{4}$ potential
Our result for the spectrum in the scaling regime can be compared with those found in previous (lower-resolution) simulations [19, 22], whereas the results of [40] have been obtained assuming a matter-dominated background. First, we notice that the width $\delta$ was not included as a parameter in [19, 40, 22]. Additionally, [19, 22] only provide fits to the high frequency region of the spectrum, rather than a fit to the full spectrum. Concerning the slope of the spectrum in this range ($f\geq f_{p}$), our results are close to the usually adopted value $\beta\approx 1$, derived in [19], and also agree with [22], whereas [40] finds a significantly steeper decrease in matter domination (with white noise initial conditions as in our case). Concerning the position of the peak, we also find reasonable agreement with the expectation $x_{p}\simeq 1$. Finally, concerning the amplitude parameter $\epsilon$, our result is within the errors of [19] and agrees very well with [22].
@@ B.1 Initial conditions
We have evolved the fields in a radiation-dominated background $a(\eta)=a_{i}+H_{i}\eta$, with $a_{i}$ and $H_{i}$ the scale factor and Hubble parameter at the initial simulation time respectively. At that time we set the homogeneous mode of the inflaton amplitude to zero. Over it, we impose a white noise spectrum of Gaussian fluctuations in momentum space below a certain cutoff $k<k_{*}$. The variance in position space is given by
@@
where we have set a tiny initial amplitude with an arbitrary small coefficient $c=10^{-10}$, since the scaling solution should not depend on it, and $k_{*}$ a ultraviolet cutoff. The procedure by which this spectrum is generated is explained in Sec. 7.1 of [30]. In our simulations we have fixed $k_{*}=2m$, but we have checked that the dynamics of the DW network are quite insensitive on the specific choice of $k_{*}$.
@@ B.4 Spectrum dependence on lattice parameters
Figure 19: Scalar field spectra for different bias sizes (represented by different coloured lines) extracted from simulations with $N=5000$, $L=90$, and second-order accurate spatial derivatives. The spectra are shown when the false vacuum fraction is $\mathcal{F}_{\rm fv}=0.25$ (solid), 0.05 (dashed), 0.02 (dotted), and 0.002 (solid thick). Note that for the largest $\eta_{\Delta V}$, we only show the spectra for the largest values of $\mathcal{F}_{\rm fv}$, since smaller values cannot be attained within the range of validity of our simulations.
B.5 Spectrum dependence on bias size
In the simulations presented in Sec. III we have fixed the size of the bias term (the coefficient $q$) such that $\eta_{\Delta V}=14$. In order to assess the robustness of our GW spectrum parametrization under changes of the bias size, in Fig. 19 we compare the final GW spectrum obtained with one realization of the case $\eta_{\Delta V}=14$, with another one for the slightly larger value $\eta_{\Delta V}=17$. We observe that both spectra are remarkably similar in the region $x<50$.
Ideally, we would like to simulate the gravitational waves for larger $\eta_{\Delta V}$, in order to further confirm the robustness of our spectrum parametrization. However, this delays the annihilation of the domain walls and hence requires larger lattices, beyond our available computational resources. Alternatively, we have been able to simulate the scalar field only (without gravitational waves) in lattices of up to $N=5000$ points per dimension, which successfully captures the annihilation process for smaller bias sizes. In Fig. 19 we compare the scalar field spectrum, i.e. $\Delta_{\phi}\equiv k^{3}\lvert\phi_{k}\lvert^{2}/(2\pi^{2})$, for different bias sizes in the range $\eta_{\Delta V}\in[14,31]$, at different times during the annihilation. More specifically, we compare the spectra when the false vacuum fraction (i.e. the fraction of the lattice volume where $\phi>0$) attains the values $\mathcal{F}_{\rm fv}=0.25,0.05,0.02,0.002$. We can observe that the shapes of the scalar spectrum at equal values of $\mathcal{F}_{\rm fv}$ are similar for different values of $\eta_{\Delta V}$, and since the scalar field sources the GW spectrum, we expect the latter to also be roughly independent of the bias size.
B.6 Comparison with previous results
@@
In Fig. 20 we compare our prediction for the shape of the GW spectrum in the biased $\phi^{4}$ case with a time-independent bias, with the one obtained in Ref. [37] for a time-dependent one (see the top panel of Fig. 3 of that work). Both spectra agree quite well in the position of the peak and the slope of the UV tail in the region $0.3\lesssim k\lesssim 5$. On the other hand, the IR tail observed in [37] is steeper than ours. Both spectra also show a plateau at high frequencies, with a difference of a factor two in its position.